Can you please help me with this fraction? Thank you.
2(3/4x + 2) = 4 - 1/2x

Found 2 solutions by Alan3354, checkley79:
Answer by Alan3354(69443)

Can you please help me with this fraction? Thank you.
2(3/4x + 2) = 4 - 1/2x
------------
Maybe.
What do you want to do with it?

Answer by checkley79(3341)

2(3/4x + 2) = 4 - 1/2x
6x/4+4=4-x/2 subtract 4.
6x/4=-x/2 cross multiply.
2*6x=-4x
12x+4x=0
16x=0
x=0 ans.
proof:
2((3*0)/4+2)=4-0/2
2*0+2*2=4-0
4=4